A 12V battery has an internal resistance of 0.05Ω. Two loads are connected in parallel across its terminal, the first drawing a current of 12A and the other taking a power of 200 watts. Determine the load terminals voltage?

1
Expert's answer
2021-11-11T02:08:01-0500

"I=E/(R+r)\to IR+0.05I=12\to"


"U+0.05I=12"


"I=I_1+I_2=12+200\/U" .So, we have


"U+(12+200/U)\cdot0.05=12"


"U^2-11.4U+10=0\to U=10.4\ (V)" . Answer